OpenStudy (anonymous):

using Euclidean algorithm find GCD(1071,462).

ganeshie8 (ganeshie8):

\(1071 = \color{red}{462}\times 2 + \color{green}{147} \)

ganeshie8 (ganeshie8):

next, check if the remainder 147 goes evenly in 462

ganeshie8 (ganeshie8):

\(462 = \color{red}{147}\times 3 + \color{green}{21} \) no luck yet, continue next iteration wid 147 and 21

OpenStudy (kc_kennylau):

Wait

OpenStudy (kc_kennylau):

I saw that before

ganeshie8 (ganeshie8):

??

OpenStudy (kc_kennylau):

nvm

OpenStudy (anonymous):

only that ? @ganeshie8

ganeshie8 (ganeshie8):

you need to continue till the remainder divides evenly or in simple words : keep going until u get a 0 remainder.

ganeshie8 (ganeshie8):

\(1071 = \color{red}{462}\times 2 + \color{green}{147} \) \(462 = \color{red}{147}\times 3 + \color{green}{21} \) \(147 = \color{red}{21} \times 7 + \color{green}{0} \) <<<<<<<< since 21 goes evenly in 147, 21 is the gcd(1071, 462)